# Do different prompting methods yield a common task representation in language models?

### v2

@@ Abstract
Demonstrations and instructions are two primary approaches for prompting language models to perform in-context learning (ICL) tasks. Do identical tasks elicited in different ways result in similar representations of the task? An improved understanding of task representation mechanisms would offer interpretability insights and may aid in steering models. We study this through function vectors (FVs), recently proposed as a mechanism to extract few-shot ICL task representations. We generalize FVs to alternative task presentations, focusing on short textual instruction prompts, and successfully extract instruction function vectors that promote zero-shot task accuracy. We find evidence that demonstration- and instruction-based function vectors leverage different model components, and offer several controls to dissociate their contributions to task performance. Our results suggest that different task promptings forms do not induce a common task representation through FVs but elicit different, partly overlapping mechanisms. Our findings offer principled support to the practice of combining instructions and task demonstrations, imply challenges in universally monitoring task inference across presentation forms, and encourage further examinations of LLM task inference mechanisms.
@@ 2.1 Generalizing function vectors beyond in-context demonstrations
[…]
@ Real texts: Intuition: sample texts from a natural corpus that convey no task information but otherwise match the task specifications. We score token sequences from a chosen corpus under the model and sample $\tilde{q}_{m}^{t}$ with approximately the same length and probability as $q_{m}^{t}$.
@ Other task specifications: Intuition: sample task specifications for other tasks $t^{\prime}\neq t$ that likely convey no information for task $t$. We score task specifications $q_{m^{\prime}}^{t^{\prime}}$ generated for other tasks $t^{\prime}\in\mathcal{T}$, and again sample ones with approximately the same length and probability as $q_{m}^{t}$.
@@
Although the method is more general, here we use zero-shot textual instructions as our $Q_{t}$. We generate a candidate set of textual instructions for each task by querying Llama-3.1-405B (Llama Team, , 2024) using a $K$-shot ICL prompt for the task, and ask it to generate 10 instruction prompts for the task it infers from the provided demonstrations (see Appendix A for the templates). We repeat this procedure 20 times for each task $t$ and deduplicate exact repetitions to arrive at a task instruction set $Q_{t}$. We repeat this procedure twice, once encouraging the model to generate short instructions and once with no such encouragement. We provide sample generated instructions of both lengths in Section A.1 and corresponding uninformative baselines (using all three approaches) in Section B.2. For our ‘real texts’ uninformative baseline, we sample texts from WikiText-103-v1 (Merity et al., , 2016).
@@ 2.2 Experimental conditions
Tasks. We consider the same set of tasks and datasets used by Todd et al., (2024). We omit a few classification datasets where successfully predicting the next token requires an understanding of the format that is facilitated by demonstrations but not necessarily by minimal textual instructions (we retain a total of 50 datasets; see Appendix D for the full list). We follow Todd et al., (2024) in computing the top sets of heads only over datasets where a model surpasses chance performance. As our approach requires 20 successful prompts for each of the five best instructions to compute the mean task-conditioned activations, we also omit tasks where a model fails to pass this number of prompts (which, in most cases, means it is also below chance accuracy and was already omitted ).
Textual instructions and uninformative baselines. We evaluate each model on each task using one random seed six separate times, for the short ($\leq 16$ tokens) vs. longer (unbounded) instructions crossed with all three uninformative baseline approaches (§2.1). We observe minimal deviation in top heads across these conditions (see Section H.3), so we average the causal effects across all six to compute the top heads. We then report final evaluation accuracies averaged over the results with both sets of mean activations, those generated with short instructions and those from longer ones.
@@
Evaluation and Comparison Logic. We structure our evaluations to facilitate comparison between FVs derived from demonstrations and instructions. We use the $|\mathcal{A}|=20$ top heads in both settings, consistent with Todd et al., (2024) for the 7B models. We follow Todd et al., (2024) in evaluating function vectors as an additive intervention to the residual stream after the $|L/3|$ layer (layer 9 for the Llama-3.2-3B models and layer 11 for the Llama-3.1-8B and OLMo-2-1123-7B ones). We report two evaluation settings: 10-shot with shuffled labels ($\tilde{p}_{i}^{t}$ in §2.1) and 0-shot (with no instructions). We focus our assessment of each FV in the setting matching its extraction: shuffled 10-shot for demonstration FVs and 0-shot for instruction FVs. In addition to the FV evaluations, we report a baseline of the evaluation setting without adding the function vector. We report accuracies using informative task presentations for each evaluation in Appendix F: instructed 0-shot (averaging over the top $J=5$ instructions for each model and task), and 10-shot (without label shuffling).
@@ 3.1 Function vectors extracted from instructions facilitate zero-shot task accuracy
Figure 2A summarizes our evaluation of demonstration function vectors (left) and instruction-based FVs (right). We find that our adaptation of the FV extraction procedure succeeds and increases zero-shot task accuracy from below 20% to above 50% in the best models. A notable exception is the base Llama-3.1-8B model, for which the $|L/3|$ intervention depth appears suboptimal (see Figure 11). We also observe that, unsurprisingly, instruction FVs are substantially more effective in post-trained models. Finally, while instruction FVs are highly effective in the zero-shot setting, they fail to match the accuracy demonstration FVs attain in the shuffled 10-shot condition, and struggle when evaluated in the shuffled 10-shot condition themselves (see Figure 12).
Figure 3: (A-D) Demonstrations and instructions elicit mostly distinct attention heads. We visualize the sets of top heads identified only with demonstrations, only with instructions, and shared in both cases. Only a few top heads are shared each in model (see Section H.1 for additional models). We observe that top instruction FV heads tend to be in earlier layers of the post-trained (-Instruct) models, compared to the base versions, and that shared heads tend to arise in later layers. (E-F) Shared head activation similarity depends on layer, but not on post-training. We plot the cosine similarity of activations in shared heads between three sets of task-conditioned activations: those elicited by demonstrations, by short instructions, and by longer instructions. We observe increasing similarity between demonstrations and instructions in later model layers, but show no effect of post-training on similarity. (G-H) ‘Heterogeneous’ function vectors show demonstration-instruction asymmetry. We construct ‘heterogeneous’ FVs using attention heads localized with instructions and demonstration task activations (left data points) to the opposite combination, using demonstration-localized heads and instruction task activations (right). Compared to the ‘homogeneous’ function vectors (middle), we mostly observe a smaller accuracy drop when using instruction FV heads to read from demonstration mean activations than vice versa.
@@
Constructing function vectors seems to benefit from model capabilities. Smaller Llama-3.2-1B models show a lower accuracy increase with instruction FVs; conversely, newer OLMo-2-1124-7B models show a greater increase (see Section G.3). This capacity is not unique to the latest models; results with Llama-2-7B models qualitatively match newer models. Finally, we evaluate the effect of choosing a fixed intervention layer in Figure 11. If we choose the optimal layer for each model and evaluation, the accuracy numbers rise as expected, but the overall patterns remain the same.
@@ 3.2 Demonstration and instruction function vectors are beneficial together
If demonstrations and instructions elicit different task representations in a language model, to what extent are their representations jointly beneficial? To study this, we intervene with both FVs, adding them to the residual stream after the same $|L/3|$ layer. Our results in Figure 2B demonstrate that adding both function vectors appears consistently beneficial, except for the base Llama-3.1-8B model (for which the choice of layer seems crucial). Surprisingly, adding both FVs after the same layer does not appear to induce interference, even though these FVs were extracted independently. In Figure 11B, we report results from the highest accuracy intervention layers in a sweep over layer pairs; adding both FVs to the same layer often proved optimal. Finally, to establish how much of this effect is simply due to the magnitude of the intervention, we report results in adding the same FV twice in Figure 15. Adding the same FV twice performs similarly to adding both FVs for some models and evaluations. This suggests that some of the benefit is due to amplifying the intervention rather than the different information carried in both FVs; we leave studying the extent of these relative effects to future work. Having found behavioral evidence that these different function vectors convey different information, we next examine to what extent they share mechanisms within the model.
@@ 3.3 Instruction and demonstration FVs mostly engage different attention heads
In Figure 3A-D, we visualize the $|\mathcal{A}|=20$ top attention heads identified in each model by the FV extraction procedure. We observe that demonstrations and instructions elicit mostly distinct sets of heads. Of the 20 top heads identified, instruction and demonstration FVs share few of them – 7 in both Llama-3.2-3B models and only 4 in the Llama-3-1.8B ones. Post-training appears to move instruction FV heads closer to demonstration FV heads. The most notable change between base and post-trained model versions is in the average layer of the instruction-only heads — from a mean several layers deeper than demonstrations in base models, to almost identical depths in post-trained ones (other model classes are similar, see Appendix H). We observe higher causal scores in demonstration top heads. In §2, we define the causal score of each head as its contribution to correctly predicting the next token. Demonstration FV top heads receive substantially higher scores than instruction ones (Table 7). This suggests that task inference from demonstrations is more localized to a small set of heads than from instructions and that instruction-based task inference is more diffuse.
We also examine the mean task-conditioned activation patterns in the top heads shared between instructions and demonstrations. We compute the cosine similarities between the patterns elicited by demonstrations vs. by shorter/longer instructions (see Figure 3E-F). Expectedly, we find higher similarities between the two instruction-driven activation patterns than between the instruction and demonstration patterns. Additionally, later layers are generally more similar. Finally, post-training appears not to adapt demonstration FV heads. We observe no difference in demonstration-instruction similarities between base and post-trained models, and equal numbers of shared heads between instructions and demonstrations across model versions. Instead, post-training appears to induce an alternative task inference mechanism, one we hope to explore in future work.
@@ 3.4 Instruction-identified attention heads are more useful for building demonstration FVs than vice versa
If demonstrations and instructions share a common task inference mechanism, we would not observe a significant difference if we constructed ‘incongruent’ function vectors using top heads identified by one with activations from the other. We borrow from neuroimaging and consider the FV extraction procedure through the lens of functional localizers, viewing head identification as localization and mean activation computation as recording (Saxe et al., , 2006; Berman et al., , 2010). We evaluate both types of ‘incongruent’ FVs, those constructed with demonstration top heads and those using instruction top heads, and summarize our results in Figure 3G-H. We observe that, as expected, accuracy with these incongruent FVs falls below the regular FVs in all cases. We note an asymmetry: using instruction-localized top heads from with demonstration mean activations seems preferable to the opposite combination (with the exception of the shuffled 10-shot evaluation on the base models). This effect is sharpened when we evaluate each FV in the layer resulting in the highest accuracy, rather than fixing to $|L/3|$ (Figure 25). To explain this, we report the causal indirect effects of the top heads from each task presentation in the opposite presentation format – that is, the scores of instruction-localized heads when using demonstrations, and vice versa (Table 8). Examining the scores, we observe that instruction-localized heads are more helpful in the demonstration setting than demonstration-localized heads are in the instruction setting. We take this as evidence that the mechanism for task inference from instruction leverages attention heads that play a minor role in demonstration task inference. In contrast, the demonstration task inference mechanism uses attention heads that are less useful for instruction task inference. This supports our earlier claim about the primacy of demonstration ICL. We conclude with two control experiments to ensure that this observed asymmetry is meaningful (and not an artifact of selecting arbitrary sets of heads). We select sets of heads that are either unrelated to both FV types, or that have the lowest causal scores (Section J.1) In both cases, accuracy falls to task baseline or below, suggesting the observed effect is meaningful.
@@ 3.5 Instruction FVs from post-trained can steer base models, arise from SFT and DPO
We conclude by examining which post-training stages contribute to the ability to extract instruction FVs using the OLMo-2-1124-7B family of models (OLMo et al., , 2024). The authors release four versions of the model: a base version, one following supervised fine-tuning (-SFT), another following preference fine-tuning (-DPO), and a final version (-Instruct). Figure 4B depicts results with these four models. All OLMo models are conducive to extracting function vectors from demonstrations, but the base model instruction FVs show lower accuracies. We observe two accuracy increases, once with the SFT model and another with the DPO model. The DPO model also offers an increase in the efficacy of instruction FVs in mismatched evaluations (see Figure 29B).

### v3

@@ Abstract
Demonstrations and instructions are two primary approaches for prompting language models to perform in-context learning (ICL) tasks. Do identical tasks elicited in different ways result in similar representations of the task? An improved understanding of task representation mechanisms would offer interpretability insights and may aid in steering models. We study this through function vectors (FVs), recently proposed as a mechanism to extract few-shot ICL task representations. We generalize FVs to alternative task presentations, focusing on short textual instruction prompts, and successfully extract instruction function vectors that promote zero-shot task accuracy. We find evidence that demonstration- and instruction-based function vectors leverage different model components, and offer several controls to dissociate their contributions to task performance. Our results suggest that different task prompting forms do not induce a common task representation through FVs but elicit different, partly overlapping mechanisms. Our findings offer principled support to the practice of combining instructions and task demonstrations, imply challenges in universally monitoring task inference across presentation forms, and encourage further examinations of LLM task inference mechanisms.
@@ 2.1 Generalizing function vectors beyond in-context demonstrations
One of our contributions is to generalize the function vector method described above from demonstrations to alternative task specifications. We denote by $Q_{t}$ a set of task specifications (of any form) for the task $t$ (e.g., the instructions in Figure 1). For a query example $(x_{iq},y_{iq})$ (Egypt $\Rightarrow$ Cairo in Figure 1), we sample $q_{m}^{t}\in Q_{t}$ and construct prompts $p_{i}^{t}$ as $p_{i}^{t}=[q_{m}^{t},x_{iq}]$ Next, we address the challenge of creating uninformative baselines $\tilde{q}_{m}^{t}$ for the task specifications $q_{m}^{t}$. We consider three approaches to generating these baselines, and see further details and examples in Appendix˜B:
• Equiprobable token sequences. Intuition: sample token sequences that are similarly likely under the model but are unrelated to the task. Each $q_{m}^{t}$ is encoded as a token sequence $q_{m}^{t}=[w_{1},w_{2},...,w_{L}]$ whose probability under the model $f$ is $P(q_{m}^{t})=\prod_{l\leq L}P(w_{l}\mid w_{<l})=f(w_{<l})[w_{l}]$. Starting from the BOS token, we sample $\tilde{w}_{l}$ to approximately match the conditional probabilities $f(\tilde{w}_{<l})[\tilde{w}_{l}]\approx f(w_{<l})[w_{l}]$, and construct $\tilde{q}_{m}^{t}=[\tilde{w}_{1},\tilde{w}_{2},\cdots,\tilde{w}_{L}]$.
• Real texts: Intuition: sample texts from a natural corpus that convey no task information but otherwise match the task specifications. We score token sequences from a chosen corpus under the model and sample $\tilde{q}_{m}^{t}$ with approximately the same length and probability as $q_{m}^{t}$.
• Other task specifications: Intuition: sample task specifications for other tasks $t^{\prime}\neq t$ that likely convey no information for task $t$. We score task specifications $q_{m^{\prime}}^{t^{\prime}}$ generated for other tasks $t^{\prime}\in\mathcal{T}$, and again sample ones with approximately the same length and probability as $q_{m}^{t}$.
@@
Although the method is more general, here we use zero-shot textual instructions as our $Q_{t}$. We generate a candidate set of textual instructions for each task by querying Llama-3.1-405B (Llama Team, , 2024) using a $K$-shot ICL prompt for the task, and ask it to generate 10 instruction prompts for the task it infers from the provided demonstrations (see Appendix˜A for the templates). We repeat this procedure 20 times for each task $t$ and deduplicate exact repetitions to arrive at a task instruction set $Q_{t}$. We repeat this procedure twice, once encouraging the model to generate short instructions and once with no such encouragement. We provide sample generated instructions of both lengths in Section˜A.1 and corresponding uninformative baselines (using all three approaches) in Section˜B.2. For our ‘real texts’ uninformative baseline, we sample texts from WikiText-103-v1 (Merity et al., , 2016).
@@ 2.2 Experimental conditions
Tasks. We consider the same set of tasks and datasets used by Todd et al., (2024). We omit a few classification datasets where successfully predicting the next token requires an understanding of the format that is facilitated by demonstrations but not necessarily by minimal textual instructions (we retain a total of 50 datasets; see Appendix˜D for the full list). We follow Todd et al., (2024) in computing the top sets of heads only over datasets where a model surpasses chance performance. As our approach requires 20 successful prompts for each of the five best instructions to compute the mean task-conditioned activations, we also omit tasks where a model fails to pass this number of prompts (which, in most cases, means it is also below chance accuracy and was already omitted ).
Textual instructions and uninformative baselines. We evaluate each model on each task using one random seed six separate times, for the short ($\leq 16$ tokens) vs. longer (unbounded) instructions crossed with all three uninformative baseline approaches (§2.1). We observe minimal deviation in top heads across these conditions (see Section˜H.3), so we average the causal effects across all six to compute the top heads. We then report final evaluation accuracies averaged over the results with both sets of mean activations, those generated with short instructions and those from longer ones.
@@
Evaluation and Comparison Logic. We structure our evaluations to facilitate comparison between FVs derived from demonstrations and instructions. We use the $|\mathcal{A}|=20$ top heads in both settings, consistent with Todd et al., (2024) for the 7B models. We follow Todd et al., (2024) in evaluating function vectors as an additive intervention to the residual stream after the $|L/3|$ layer (layer 9 for the Llama-3.2-3B models and layer 11 for the Llama-3.1-8B and OLMo-2-1123-7B ones). We report two evaluation settings: 10-shot with shuffled labels ($\tilde{p}_{i}^{t}$ in §2.1) and 0-shot (with no instructions). We focus our assessment of each FV in the setting matching its extraction: shuffled 10-shot for demonstration FVs and 0-shot for instruction FVs. In addition to the FV evaluations, we report a baseline of the evaluation setting without adding the function vector. We report accuracies using informative task presentations for each evaluation in Appendix˜F: instructed 0-shot (averaging over the top $J=5$ instructions for each model and task), and 10-shot (without label shuffling).
@@ 3.1 Function vectors extracted from instructions facilitate zero-shot task accuracy
Figure 2A summarizes our evaluation of demonstration function vectors (left) and instruction-based FVs (right). We find that our adaptation of the FV extraction procedure succeeds and increases zero-shot task accuracy from below 20% to above 50% in the best models. A notable exception is the Llama-3.1-8B model, for which the $|L/3|$ intervention depth appears suboptimal (see results from other intervention layers in Figure 11). On the other hand, the Llama-3.2 models may benefit from being distilled from larger, post-trained Llama-3.1 models. The top $|\mathcal{A}|=20$ attention heads also represent a slightly larger fraction of attention heads in the smaller models (about $3\%$ in the 3B models vs. $2\%$ in the 8B ones). We also observe that, unsurprisingly, instruction FVs are substantially more effective in post-trained models. Finally, while instruction FVs are highly effective in the zero-shot setting, they fail to match the accuracy demonstration FVs attain in the shuffled 10-shot condition, and struggle when evaluated in the shuffled 10-shot condition themselves (see Figure 12). We replicate findings 1 and 2 (below) with other models: see Figures 13-15, in appendices G.3-G.4.
Figure 3: (A-D) Demonstrations and instructions elicit mostly distinct attention heads. We visualize the sets of top heads identified only with demonstrations, only with instructions, and shared in both cases. Only a few top heads are shared each in model (see Section˜H.1 for additional models). We observe that top instruction FV heads tend to be in earlier layers of the post-trained (-Instruct) models, compared to the base versions, and that shared heads tend to arise in later layers. (E-F) Shared head activation similarity depends on layer, but not on post-training. We plot the cosine similarity of activations in shared heads between three sets of task-conditioned activations: those elicited by demonstrations, by short instructions, and by longer instructions. We observe increasing similarity between demonstrations and instructions in later model layers, but show no effect of post-training on similarity. (G-H) ‘Heterogeneous’ function vectors show demonstration-instruction asymmetry. We construct ‘heterogeneous’ FVs using attention heads localized with instructions and demonstration task activations (left data points) to the opposite combination, using demonstration-localized heads and instruction task activations (right). Compared to the ‘homogeneous’ function vectors (middle), we mostly observe a smaller accuracy drop when using instruction FV heads to read from demonstration mean activations than vice versa.
@@
Constructing function vectors seems to benefit from model capabilities. Smaller Llama-3.2-1B models show a lower accuracy increase with instruction FVs; conversely, newer OLMo-2-1124-7B models show a greater increase (see Section˜G.4). This capacity is not unique to the latest models; results with Llama-2-7B models qualitatively match newer models. Finally, we evaluate the effect of choosing a fixed intervention layer in Figure 11. If we choose the optimal layer for each model and evaluation, the accuracy numbers rise as expected, but the overall patterns remain the same.
@@ 3.2 Demonstration and instruction function vectors are beneficial together
If demonstrations and instructions elicit different task representations in a language model, to what extent are their representations jointly beneficial? To study this, we intervene with both FVs, adding them to the residual stream after the same $|L/3|$ layer. Our results in Figure 2B demonstrate that adding both function vectors appears consistently beneficial, except for the base Llama-3.1-8B model (for which the choice of layer seems crucial). Surprisingly, adding both FVs after the same layer does not appear to induce interference, even though these FVs were extracted independently. In Figure 11B, we report results from the highest accuracy intervention layers in a sweep over layer pairs; adding both FVs to the same layer often proved optimal. Finally, to establish how much of this effect is simply due to amplifying the intervention, we report the results of adding the same FV twice in Figure 18. For some models and evaluations, adding the same FV twice performs similarly to adding both FVs. This suggests that some of the benefit is due to the magnitude of the intervention rather than the different information carried in both FVs; we leave studying the extent of these relative effects to future work. Having found behavioral evidence that these different function vectors convey different information, we next examine to what extent they share mechanisms within the model.
@@ 3.3 Instruction and demonstration FVs mostly engage different attention heads
In Figure 3A-D, we visualize the $|\mathcal{A}|=20$ top attention heads identified in each model by the FV extraction procedure. We observe that demonstrations and instructions elicit mostly distinct sets of heads. Of the 20 top heads identified, instruction and demonstration FVs share few of them – 7 in both Llama-3.2-3B models and only 4 in the Llama-3-1.8B ones. Post-training appears to move instruction FV heads closer to demonstration FV heads. The most notable change between base and post-trained model versions is in the average layer of the instruction-only heads — from a mean several layers deeper than demonstrations in base models, to almost identical depths in post-trained ones (other model classes are similar, see Appendix˜H). We observe higher causal scores in demonstration top heads. In §2, we define the causal score of each head as its contribution to correctly predicting the next token. Demonstration FV top heads receive substantially higher scores than instruction ones (Table 7). This suggests that task inference from demonstrations is more localized to a small set of heads than from instructions and that instruction-based task inference is more diffuse.
We also examine the mean task-conditioned activation patterns in the top heads shared between instructions and demonstrations. We compute the cosine similarities between the patterns elicited by demonstrations vs. by shorter/longer instructions (see Figure 3E-F). Expectedly, we find higher similarities between the two instruction-driven activation patterns than between the instruction and demonstration patterns. Additionally, later layers are generally more similar. Finally, post-training appears not to adapt demonstration FV heads. We observe no difference in demonstration-instruction similarities between base and post-trained models, and equal numbers of shared heads between instructions and demonstrations across model versions. Instead, post-training appears to induce a different, separate mechanism for instruction task inference, one we hope to explore in future work.
@@ 3.4 Instruction-identified attention heads are more useful for building demonstration FVs than vice versa
If demonstrations and instructions share a common task inference mechanism, we would not observe a significant difference if we constructed ‘incongruent’ function vectors using top heads identified by one with activations from the other. We borrow from neuroimaging and consider the FV extraction procedure through the lens of functional localizers, viewing head identification as localization and mean activation computation as recording (Saxe et al., , 2006; Berman et al., , 2010). We evaluate both types of ‘incongruent’ FVs, those constructed with demonstration top heads and those using instruction top heads, and summarize our results in Figure 3G-H. We observe that, as expected, accuracy with these incongruent FVs falls below the regular FVs in all cases. We note an asymmetry: using instruction-localized top heads from with demonstration mean activations seems preferable to the opposite combination (with the exception of the shuffled 10-shot evaluation on the base models). This effect is sharpened when we evaluate each FV in the layer resulting in the highest accuracy, rather than fixing to $|L/3|$ (Figure 28). To explain this, we report the causal indirect effects of the top heads from each task presentation in the opposite presentation format – that is, the scores of instruction-localized heads when using demonstrations, and vice versa (Table 8). Examining the scores, we observe that instruction-localized heads are more helpful in the demonstration setting than demonstration-localized heads are in the instruction setting. We take this as evidence that the mechanism for task inference from instruction leverages attention heads that play a minor role in demonstration task inference. The development of such a mechanism may also explain the depth alignment shifts we previously observed between base and post-trained models (subsection 3.3, Figure 3A-D) In contrast, the demonstration task inference mechanism uses attention heads that are less useful for instruction task inference. This supports our earlier claim about the primacy of demonstration ICL. We conclude with two control experiments to ensure that this observed asymmetry is meaningful (and not an artifact of selecting arbitrary sets of heads). We select sets of heads that are either unrelated to both FV types, or that have the lowest causal scores (Section˜J.1) In both cases, accuracy falls to task baseline or below, suggesting the observed effect is meaningful.
@@ 3.5 Instruction FVs from post-trained can steer base models, arise from SFT and DPO
We conclude by examining which post-training stages contribute to the ability to extract instruction FVs using the OLMo-2-1124-7B family of models (OLMo et al., , 2024). The authors release four versions of the model: a base version, one following supervised fine-tuning (-SFT), another following preference fine-tuning (-DPO), and a final version (-Instruct). Figure 4B depicts results with these four models. All OLMo models are conducive to extracting function vectors from demonstrations, but the base model instruction FVs show lower accuracies. We observe two accuracy increases, once with the SFT model and another with the DPO model. The DPO model also offers an increase in the efficacy of instruction FVs in mismatched evaluations (see Figure 33B). Finally, we note that post-training appears to have no impact on demonstration-based ICL, as evidenced by the lack of variability in Figure 4B (left).